OpenStudy (anonymous):

what is the equation of the line (-2,1) and (0,1)

OpenStudy (anonymous):

find the slop first m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

Hint: Find Slope. \[m =\frac{ y2 - y1 }{ x2 - x1 }\] Put into Point-Slope Form. *Substitute the x and y values of a give point into the equation. **Remember m is slope. \[y - y1 = m(x - x1)\]

OpenStudy (anonymous):

Once you have put everything into Slope-Intercept form; Distribute. Solve for y.